OpenStudy (smurfy14):

write the equation for a line that passes through (-2,3) and is perpendicular to the line 3x+4y=-7. Write in point-slope form.

OpenStudy (anonymous):

3x+4y=-7 4y=-3x-7 divide by 4 all through y=-3/4x-7/4 gradient of this equation is -3/4 (-2,3) perpendicular is normally equal to -1 when you multiply both the gradients of the two that is; -3/4 * x = -1 x=4/3 this tends to be the gradient for the second equation to find the equation for (-2,3) and we know the gradient is 4/3 y=mx+c y=4/3x + c (-2,3) is represnted by (x,y) find change in y / change in x y+3/x+2 = 4/3 3(y+3)=4(x+2) 3y+9=4x+8 3y=4x+8-9 3y=4x-1

OpenStudy (anonymous):

3y-4x=-1

OpenStudy (anonymous):

-4x+3y=-1